# A Community-driven vision for a new Knowledge Resource for AI

### v2

@@ 1 Introduction
A workshop at the 2025 Conference for the Advancement of Artificial Intelligence (AAAI) gathered over 50 researchers, including the authors of this paper, to explore these questions. This paper is an attempt to synthesize our discussions, and present a community view on the requirements and approach for creating a new knowledge infrastructure. We concluded that building an open engineering infrastructure for disseminating and using knowledge modules in practical applications, leveraging steady advances in knowledge representation and reasoning over the last 40 years, is a promising path forward. This includes establishing sets of conventions and social structures that should be adopted by contributors.
We begin by envisioning the type of knowledge resource required in the current context of AI, and then exploring aspects of creating it, including foundational knowledge, domain-specific knowledge, automated reasoning, and evaluation. We also review the current state of education in knowledge representation and reasoning. We conclude by presenting a community view on productive steps towards the creation of this much needed knowledge resource.
@@ 2.1 Problems solvable by a Knowledge Resource
Data-driven learning is central to modern artificial intelligence (AI). But in some cases, curated knowledge can be better. For example, in basic arithmetic, people easily outcompete chatbots such as ChatGPT because they learn how to add and multiply numbers using rules rather than by looking at a large number of examples [14].
@@
A set of concrete examples where a knowledge resource can supplement the knowledge in large language models has been put out by Wolfram|Alpha [82]. These include “distance between Chicago and Tokyo”, “3 to the power 73”, “circumference of an ellipse with axes 4 and 12”, etc.
@@ 2.2 Limits of the current practice
Knowledge resources such as Cyc [49], Wolfram|Alpha [38], Component Library [4] or its variant CoreALMLib [39] attempt to bridge this gap. Cyc, for example, targets formalization of complex knowledge patterns across the full spectrum of the human world. Because of the restricted intellectual property of Cyc and Wolfram|Alpha, limited access to them is not conducive to fostering an open-source community. OpenCyc, which is the open-source version of Cyc, is useful for limited purposes as it provides access to only a small fraction of the full knowledge base.
@@ 7.3 Evaluation in virtual environments
Virtual environments, simulated worlds and games can be effective in evaluating targeted aspects of reasoning. For example, the angry birds competition [65] has been effective at exploring reasoning about actions and qualitative reasoning. Game environments and cognitive robotics tasks can also constrain the allowed moves in ways that can explicitly force any successful program to reason.
@@ 8 Next steps towards creating a knowledge resource
There was an overwhelming consensus at the workshop that a meticulously curated knowledge resource along with the necessary tools and methodologies for its effective use is sorely needed.
@@
While some argued for intrinsic value of a knowledge resources on its own. Standalone knowledge resources already exist, for example, Common Logic Ontology Repository [33], Standard Upper Merged Ontology [58] and Basic Formal Ontologies [3]. These prior existing knowledge resources differ from the Hugging Face model, and the new knowledge resource envisioned here, in that they do not target any specific end-user task. In addition to focusing on specific tasks, the success of Hugging Face can be attributed to providing interoperability between a few prevalent deep learning models, creating a unified API that simplifies training across different model architectures and tasks, and providing a hub for model training and discovery.
@@ 9 Conclusion
Much work needs to be done to turn the community vision articulated here into an actionable research program. In addition to the initial steps suggested in the previous paragraph, the importance of research in all aspects of creating the knowledge resource is equally important. Significant research is needed in methodologies for creating both foundational and domain-specific knowledge, in effective reasoning techniques that scale and take context into account, and in translational techniques that can automatically translate a reasoning task framed in one formalism into another formalism. Enabling seamless distributed development, especially by domain experts is essential for fostering an effective community.

### v3

@@ 1 Introduction
A workshop at the 2025 Conference for the Advancement of Artificial Intelligence (AAAI) gathered over 50 researchers, including the authors of this paper, to explore these questions [80]. This paper is an attempt to synthesize our discussions, and present a community view on the requirements and approach for creating a new knowledge infrastructure. We concluded that building an open engineering infrastructure for disseminating and using knowledge modules in practical applications, leveraging steady advances in knowledge representation and reasoning over the last 40 years, is a promising path forward. This includes establishing sets of conventions and social structures that should be adopted by contributors.
We begin by envisioning the type of knowledge resource required in the current context of artificial intelligence (AI), and then exploring aspects of creating it, including foundational knowledge, domain-specific knowledge, automated reasoning, and evaluation. We also review the current state of education in knowledge representation and reasoning. We conclude by presenting a community view on productive steps towards the creation of this much needed knowledge resource.
@@ 2.1 Problems solvable by a Knowledge Resource
Data-driven learning is central to modern AI. But in some cases, curated knowledge can be better. For example, in basic arithmetic, people easily outcompete chatbots such as ChatGPT because they learn how to add and multiply numbers using rules rather than by looking at a large number of examples [15].
@@
It has been shown through the qualitative spatial reasoning benchmark, Room Space 100, that When the number of objects in a scene (n) increases from 3 to 6, GPT-4’s accuracy generally declines, for example, for n=3, GPT-4 has an accuracy of 0.55, and for n=6, an accuracy of 0.15 [54]. It has also been found that GPT-4 could solve some action-based complex puzzles – performing better when generating Python code – but an axiomatic reasoner using a proper action description language solved all puzzles correctly [41]. A set of concrete examples where a knowledge resource can supplement the knowledge in large language models has been put out by Wolfram|Alpha [90]. These include “distance between Chicago and Tokyo”, “3 to the power 73”, “circumference of an ellipse with axes 4 and 12”, etc.
@@ 2.2 Limits of the current practice
Knowledge resources such as Cyc [52], Wolfram|Alpha [39], Component Library [4] or its variant CoreALMLib [40] attempt to bridge this gap. Cyc, for example, targets formalization of complex knowledge patterns across the full spectrum of the human world. Because of the restricted intellectual property of Cyc and Wolfram|Alpha, limited access to them is not conducive to fostering an open-source community. OpenCyc, which is the open-source version of Cyc, is useful for limited purposes as it provides access to only a small fraction of the full knowledge base [64].
Demonstration of inference gaps in LLMs for spatial reasoning and action reasoning cited in the previous section have been limited to academic settings. Establishing that such inference gaps are critical for real-world applications is an open research problem.
@@ 7.3 Evaluation in virtual environments
Virtual environments, simulated worlds and games can be effective in evaluating targeted aspects of reasoning. For example, the angry birds competition [70] has been effective at exploring reasoning about actions and qualitative reasoning. Game environments and cognitive robotics tasks can also constrain the allowed moves in ways that can explicitly force any successful program to reason. Recently proposed Gardner test is an example of such a test that situates the rich tradition of General Game Playing competitions in the context of modern generative AI systems [13].
@@ 8 Next steps towards creating a knowledge resource
The AAAI workshop [80] that is the basis of the present article is part of a larger initiative [75] to formulate and execute a research program to develop a new knowledge resource for AI. The present workshop is the first step towards finding the requirements and building a community. We have already planned three additional workshops that are focused on use cases for the envisioned knowledge resource. After the AAAI workshop, a subset of the participants formulated a project to develop a proof of concept for the utility of the knowledge resource. With the wider community feedback, and the results from the proof-of-concept project, we will be in a better place to design the engineering framework and formulate the associated research problems that must be solved. For the rest of the section, we discuss the recommendations from the workshop participants, plan for the follow on workshops, and the proof of concept feasibility study.
8.1 Recommendations from the Workshop Participants
The workshop participants agreed that a meticulously curated knowledge resource along with the necessary tools and methodologies for its effective use is sorely needed.
@@ 8.1 Recommendations from the Workshop Participants
While some argued for intrinsic value of a knowledge resources on its own, standalone knowledge resources already exist, for example, Common Logic Ontology Repository [34], Standard Upper Merged Ontology [62] and Basic Formal Ontologies [3]. These prior existing knowledge resources differ from the Hugging Face model, and the new knowledge resource envisioned here, in that they do not target any specific end-user task. In addition to focusing on specific tasks, the success of Hugging Face can be attributed to providing interoperability between a few prevalent deep learning models, creating a unified API that simplifies training across different model architectures and tasks, and providing a hub for model training and discovery.
@@
8.2 Follow on Workshops
We are planning three follow on workshops that are focused on specific use cases that can benefit from a knowledge resource: education, supply chain and computational law. These topics were chosen because there exists prior work to justify further exploration. We briefly describe each of these workshops.
The workshop on education is aimed at at identifying problems that cannot be solved using LLMs alone and require knowledge representation to be created. Examples of such problems include skill graphs, precision knowledge tracing, and grounding AI in verifiable knowledge. The goal of this workshop is to formulate a large community-driven knowledge graph that would benefit education.
The workshop on supply chain is aimed at addressing critical vulnerabilities in the global supply chain exposed by the pandemic and infrastructure failures. We will explore several potential supply chain domains (for example, minerals, plastics, water supply, etc.) with an eye towards identifying a domain where the data is easily available, where a global view of the supply chain could be enabled by creating a rich semantic model.
The workshop on computational law is aimed at creating a community of users for a national library of laws that are represented as computer code. Preliminary work exists in codifying local building codes suggesting that scaling it to a national level will address significant inefficiency in regulatory compliance landscape.
8.3 Proof of Concept for a Knowledge Resource
To adapt the Hugging Face model to knowledge modules, the major roadblock is not the design of a portal itself but to identify concrete problems and use cases where the use of a knowledge resource makes a significant difference. Dr Alessandro Oltramari, the president of the Carnegie Bosch Institute, who attended the workshop, came forward with a set of use cases from Bosch that could be used to establish the value of a knowledge resource. Consequently, Dr. Chaudhri and Prof. Shimizu worked with Bosch to define three use cases: inference gaps in LLMs, robotic skill learning and root cause analysis. We briefly explain each of these use cases.
Through a qualitative reasoning benchmark, Room Space 100, it has been shown that the accuracy of GPT-4 declines from 0.55 to 0.15 when the number of objects in a scene increase from 3 to 6 [54]. In contrast a qualitative spatial reasoner for the same task is complete and correct regardless of the number of users. We will evaluate this claim in the context of a robotic planning use case case provided by Bosch. Bosch currently uses LLMs for both high-level and low-level planning [73]. In our experiment, we will replace the LLM used by high-level planner by a symbolic planner that uses a spatial reasoner, and compare their performance.
Robotic skill learning can benefit by leveraging a knowledge resource. Envision a scenario in which a robot learns to Pick up a hard object, Orient it for insertion, and then finally performs the Insert. In the current approach, a robot would learn this process as one unit. By using a knowledge resource, it could map the process to individual steps such as Pick, Orient, etc. which are more generalizable. For example, if a robot has learned to pick up a hard object while doing spark plug insertion, it could use the same skill while moving a hard object from one place to another. Starting from the skill library provided by Bosch [73], we will evaluate how many of these could be mapped to existing resources to enable generalizable robotic skill learning.
Root cause analysis in manufacturing involves identifying the cause for a defect. For example, temperature spike could be caused by machine vibration which could be caused by an alignment error. In the current approach, the root cause analysis primarily relies on a machine learning algorithm that processes sensor data. If some of the causes are modeled in structured knowledge [42], root causes analysis need not rely solely only on sensor data. We will evaluate how the use of a knowledge resource improves the root cause prediction accuracy.
If the results from any of the above evaluations are positive, that will serve as a template to replicate similar applications to other use cases. Such use cases can also be the starting point for the envisioned open engineering framework.
@@ 9 Conclusion
It is natural to question the value of a curated knowledge resource in the context of modern AI. The world is captivated by ever-larger generative AI models that produce fluent text, lifelike images, and powerful predictions. These systems dominate the headlines, attract billions in investment, and fuel the race for AI supremacy. No matter how impressive, the generative AI systems have dangerous flaws: they reflect the data they were trained on, lack formal guarantees, and will always have inference gaps. They offer performance without understanding, fluency without truth. And yet, the incentives of today’s AI ecosystem—benchmarks, funding, and hype—reward more scale, not more sense.
That is why the kinds of knowledge resources pioneered in the classical AI must return. Knowledge resources built on logic, rules, and meaning offer exactly what generative AI lacks: transparency, justification, formal guarantees. They let us encode shared human knowledge and values, instead of outsourcing everything to inscrutable models. Alone, symbolic AI once faltered. But in partnership with deep learning, it can give us systems that are both powerful and trustworthy.
The research culture of deep learning has much to teach us. Abundant tools, open-source frameworks and shared data sets that are the staple of deep learning research make it easy for newcomers to experiment. Symbolic AI knowledge resources lack comparable support and require painstaking modeling and domain expertise. This creates barriers to their adoption by AI engineers.
Our hope through the present workshop, and the follow on activities is to re-energize the knowledge representation and reasoning community in creating knowledge resources that parallel the deep learning models by adapting the engineering practices exemplified by Hugging Face. We must, however, start small and demonstrate the value and effectiveness of knowledge resources within the modern context of Generative AI systems. We have embarked on that journey by identifying industry use cases where LLMs have inference gaps, machine learning is done on very specific skills, and causal reasoning is drowning in data. We will evaluate the usefulness of existing knowledge resources in these three contexts, and prototype how the knowledge artifacts can be disseminated for widespread use. We hope that this model of working from uses cases to delivering easily reusable engineering products will inspire others in the knowledge representation and reasoning community to undertake similar efforts in their own spheres.
Much work needs to be done to define actionable research program. Based on the results of the proof of concept study mentioned in the previous section, we hope to define research projects in all aspects of creating the knowledge resource: methodologies for creating both foundational and domain-specific knowledge, effective reasoning techniques that scale and take context into account, and in translational techniques that can automatically translate a reasoning task framed in one formalism into another formalism. Enabling seamless distributed development, especially by domain experts is essential for fostering an effective community. Last but not the least, we must pay attention to considering how to involve contributors from diverse geographic, linguistic, cultural, and institutional backgrounds, including underrepresented or low-resource communities.